# Bounding causal effects with an unknown mixture of informative and non-informative missingness

### v3

@@ 4 Estimation and inference
In practice, since the nuisance functions $(e,\pi_{0},\pi_{1},\mu_{0},\mu_{1})$ are unknown, we cannot compute the bounds exactly. Instead, we must estimate the bounds as functions of these quantities using observed data. To do this, we propose using an influence-function-based estimation approach. Estimators based on influence functions typically have a so-called “doubly-robust” property, enabling them to achieve parametric rates of convergence while still allowing for the use of non-parametric and machine learning estimation methods for the nuisance functions (e.g. the propensity score $e$, the missingness probabilities $\pi_{a}$, and the outcomes models $\mu_{a}$ – collectively denoted by $\eta=\{e(\cdot),\pi_{0}(\cdot),\pi_{1}(\cdot),\mu_{0}(\cdot),\mu_{1}(\cdot)\}$). This property occurs because the errors of influence-function-based estimators are products of the errors of nuisance function estimates (call them ${\widehat{\eta}}=\{{\widehat{e}}(\cdot),{\widehat{\pi}}_{0}(\cdot),{\widehat{\pi}}_{1}(\cdot),{\widehat{\mu}}_{0}(\cdot),{\widehat{\mu}}_{1}(\cdot)\}$). Consequently, the overall estimator can achieve a faster rate of convergence even when the individual components of $^{\widehat{\eta}}$ converge at slower rates (see kennedy2022semiparametric for more details). As a result, we require minimal assumptions on the data distribution to estimate the bounds.
All proposed bounds (and the point identification) can be expressed as linear combinations of elements of the vector $\boldsymbol{\theta}=\mathbb{E}\{\zeta(\mathbf{X})\},$ where $\zeta(\mathbf{x})=\{\mu_{0}(\mathbf{x}),\mu_{1}(\mathbf{x}),\pi_{0}(\mathbf{x}),\pi_{1}(\mathbf{x}),\mu_{0}(\mathbf{x})\pi_{0}(\mathbf{x}),\mu_{1}(\mathbf{x})\pi_{1}(\mathbf{x})\}$. Because the influence function of a linear combination of functionals is a linear combination of the same influence functions (kennedy2022semiparametric), it suffices to state the influence function of the components of $\mathbb{E}\{\zeta(\mathbf{X})\}$. All bounding quantities can be estimated based on the following influence functions:
@@
The uncentered influence functions for the functionals $\mathbb{E}\{\mu_{a}(\mathbf{X})\}$, $\mathbb{E}\{\pi_{a}(\mathbf{X})\}$, and $\mathbb{E}\{\mu_{a}(\mathbf{X})\pi_{a}(\mathbf{X})\}$ are given by $\varphi_{1,a}({\bf O};\eta_{a}),\varphi_{2,a}({\bf O};\eta_{a})$, $\varphi_{3,a}({\bf O};\eta_{a})$, respectively, where
@@
$\varphi_{3,a}({\bf O};\eta)$ $=\varphi_{1,a}({\bf O};\eta_{a})\pi_{a}(\mathbf{X})+\varphi_{2,a}({\bf O};\eta_{a})\mu_{a}(\mathbf{X})-\mu_{a}(\mathbf{X})\pi_{a}(\mathbf{X}).$
@@
Suppose the target of inference is $\theta_{\mbox{$a$}}={\mbox{$a$}}^{{\sf\scriptscriptstyle{T}}}\boldsymbol{\theta}$, a linear combination of $\boldsymbol{\theta}$. Collect the influence functions into the vector $\mathbf{\varphi}({\bf O},\eta)=\left\{\varphi_{1,0}({\bf O},\eta),\varphi_{1,1}({\bf O},\eta),\varphi_{2,0}({\bf O},\eta),\varphi_{2,1}({\bf O},\eta),\varphi_{3,0}({\bf O},\eta),\varphi_{3,1}({\bf O},\eta)\right\}.$ All the quantities discussed above, including the bounds, can be written in this form. Corollary 2 in Appendix 9 provides explicit expressions for all previously considered quantities. We now describe the asymptotic behavior of the estimators of these quantities.
@@
${\widehat{\theta}}_{\mbox{$a$}}-\theta_{\mbox{$a$}}$ $=\mathbb{P}_{n}[{\mbox{$a$}}^{{\sf\scriptscriptstyle{T}}}\varphi({\bf O};\eta)-\theta_{\mbox{$a$}}]+o_{p}(n^{-1/2})$
$+\mathcal{O}_{p}(\sum_{a=0,1}\|\mu_{a}(\mathbf{X})-\hat{\mu}_{a}(\mathbf{X})\|\left\{\|\pi_{a}(\mathbf{X})-\hat{\pi}_{a}(\mathbf{X})\|+\|e_{a}(\mathbf{X})-\hat{e}_{a}(\mathbf{X})\|\right\}$
$+\|e_{a}(\mathbf{X})-\hat{e}_{a}(\mathbf{X})\|\|\pi_{a}(\mathbf{X})-\hat{\pi}_{a}(\mathbf{X})\|).$
Therefore, if
$\omega(\mathbf{X})$ $=\sum_{a=0,1}\|\mu_{a}(\mathbf{X})-\hat{\mu}_{a}(\mathbf{X})\|\left\{\|\pi_{a}(\mathbf{X})-\hat{\pi}_{a}(\mathbf{X})\|+\|e_{a}(\mathbf{X})-\hat{e}_{a}(\mathbf{X})\|\right\}$
$+\|e_{a}(\mathbf{X})-\hat{e}_{a}(\mathbf{X})\|\|\pi_{a}(\mathbf{X})-\hat{\pi}_{a}(\mathbf{X})\|=o_{p}(n^{-1/2})$
it follows by the Central Limit Theorem that
$\sqrt{n}[{\widehat{\theta}}_{\mbox{$a$}}-\theta_{\mbox{$a$}}]\stackrel{{\scriptstyle d}}{{\to}}\mathcal{N}(0,{\mbox{$a$}}^{{\sf\scriptscriptstyle{T}}}\mathbb{V}[\varphi({\bf O};\eta)]{\mbox{$a$}}).$
This theorem states that any linear combination of $\boldsymbol{\theta}$ has a limiting normal distribution with asymptotic variance given by the variance of its influence function. A valid $(1-\alpha)\times 100\%$ confidence interval may thus be constructed as
${\widehat{\theta}}_{\mbox{$a$}}\pm z_{1-\alpha/2}\sqrt{{\mbox{$a$}}^{{\sf\scriptscriptstyle{T}}}\widehat{\mathbb{V}}[\varphi({\bf O};\eta)]{\mbox{$a$}}},$
where $z_{q}$ is the $q$th quantile of the standard normal distribution. This limiting distribution holds as long as the nuisance parameters are estimated at rates that are not too slow (one way to ensure this is that all nuisance parameters converge at the $o_{p}(n^{-1/4})$ rate). Most off-the-shelf machine learning algorithms can achieve these rates under minimal assumptions on the data-generating mechanism.
An immediate consequence of this result, for example, is that the point identification in Proposition 5 corresponds to choosing ${\mbox{$a$}}=\{-1,1,0,0,(1-\tau)\delta_{0},(\tau-1)\delta_{1}\}$ such that $\theta_{\mbox{$a$}}=\Psi_{0}$ under Assumptions 1-5, 7.f, and 6.a, and an estimator for this quantity can be constructed as
@@
with corresponding confidence interval ${\widehat{\Psi}}_{0}\pm z_{1-\alpha/2}\sqrt{{\mbox{$a$}}^{{\sf\scriptscriptstyle{T}}}\widehat{\mathbb{V}}[\varphi({\bf O};{\widehat{\eta}})]{\mbox{$a$}}}$.
@@
The second-order remainder term $\omega(\mathbf{X})$ given in Theorem 1 may involve more error products than are strictly required for specific bounds. The precise second-order remainder terms for individual estimators of each bound are given in Corollary 2 in Appendix 9.
@@
Assuming $^{\widehat{\eta}}$ comes from an independent sample facilitates the statement of Theorem 1 but is not required, provided that cross-fitting is used to separate the estimation of nuisance functions from their evaluation in the influence function (see chernozhukov2018double).
@@ 5 Alternative target estimands
In some cases, however, this composite outcome may not be of interest. Building on ideas from the mediation literature (robins2022interventionist), stensrud2022separable proposed a separable direct effect (SDE) that essentially isolates the effect of the treatment on $Y$. Specifically, the SDE conceives of the treatment indicator $A$ as being comprised of two separate treatments, $A_{y}$ and $A_{d}$, where $A_{y}$ only affects the outcome and $A_{d}$ only affects $U_{I}$. That is, we can conceive of a hypothetical trial where $A_{y}$ and $A_{d}$ are jointly randomized and recover causal contrasts involving the potential outcomes $Y(a_{y},a_{d})$ for different realizations of $a_{y},a_{d}$. Here we are interested in the part of $A$ that affects $Y$ (i.e., $A_{y}$), while holding the part of $A$ that affects $U_{I}$ constant at $A_{d}=0$. We therefore first assume that $Y(a)=Y(a_{y},a_{d})$, and that $U_{I}(a)=U_{I}(a_{y},a_{d})$, and define the SDE as $\Psi_{2}=\mathbb{E}\{Y(1,0)-Y(0,0)\}$. Again, this estimand represents the causal effect via the pathways from $A$ that affect $Y$ alone ($A\boldsymbol{\rightarrow}A_{y}\rightarrow Y$, where the bold arrow indicates a deterministic relationship), leaving the pathways that affect $U_{I}$ unchanged ($A\boldsymbol{\rightarrow}A_{d}\to U_{I}\to Y$). Neither $\Psi_{1}$ or $\Psi_{2}$ are identified under the basic assumptions (Assumptions 1-5). However, these assumptions are sufficient to rewrite $\Psi_{1}$ similarly to how $\Psi_{0}$ was rewritten in Proposition 1.
@@
To identify $\Psi_{2}$, we further require the so-called “dismissible component conditions” (stensrud2022separable).
@@
This assumption is required because we do not observe the hypothetical trial that jointly randomizes $(A_{y},A_{d})$. Instead, our observed data contain only the composite treatment indicator $A$, so that $A_{y}=A_{d}$ for all units. This assumption also intuitively formalizes the idea that $A_{d}$ does not affect $Y(a_{y},a_{d})$ except via its effect on $U_{I}$, and similarly, that $A_{y}$ does not affect $U_{I}$. We refer to stensrud2022separable for more details.
@@
Unlike $\Psi_{0}$, these estimands do not depend on the unobserved quantity $\mu_{a}^{\star}(\mathbf{X})$, and therefore Assumptions 7.d-7.e are not required to bound these quantities. Instead, $\Psi_{1}$ only depends on the unobserved quantities $\pi^{\star}_{a}(\mathbf{X})$ for $a=0,1$, and $\Psi_{2}$ only on $\pi^{\star}_{0}(\mathbf{X})$. In the next section, we provide bounds for these alternative estimands analogous to the bounds in Propositions 4 and 5.
@@ 5.1 Bounds
$\ell_{2}=\tilde{\Psi}+\mathbb{E}\left[\delta_{u0}\pi_{0}(\mathbf{X})(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))\mathbb{I}\{\mu_{1}(\mathbf{X})\leq\mu_{0}(\mathbf{X})\}\right],$
$u_{2}=\tilde{\Psi}+\mathbb{E}\left[\delta_{u0}\pi_{0}(\mathbf{X})(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))\mathbb{I}\{\mu_{1}(\mathbf{X})>\mu_{0}(\mathbf{X})\}\right].$
To illustrate, consider again the running example introduced in Section 3. In this case, $\Psi_{1}=0.17$ and $\Psi_{2}=0.12$. Setting $\delta_{1u}=\delta_{0u}=1$ and $\delta_{1\ell}=\delta_{0\ell}=0$ (which hold by definition), we derive the bounds $[0.004,0.314]$ for $\Psi_{1}$ and $[0.114,0.135]$ for $\Psi_{2}$. These can be further refined using Assumption 6.b.
@@
As noted previously, Assumption 6.a is quite strong, so it is most useful as part of sensitivity analyses, as discussed in Section $3.$ Moreover, for $\Psi_{1}$, once we are simply obtaining point estimates under a range of $\delta_{1},\delta_{0}$, this is essentially equivalent to obtaining bounds $\ell_{1},u_{1}$ under different choices of $\delta_{ua},\delta_{\ell a}$.

### v4

@@ 4 Estimation and inference
In practice, since the nuisance functions $(e,\pi_{0},\pi_{1},\mu_{0},\mu_{1})$ are unknown, we cannot compute the bounds exactly. Instead, we must estimate the bounds as functions of these quantities using observed data. To do this, we propose using an influence-function-based estimation approach. Estimators based on influence functions typically have a so-called “doubly-robust” property, enabling them to achieve parametric rates of convergence while still allowing for the use of non-parametric and machine learning estimation methods for the nuisance functions (e.g. the propensity score $e$, the missingness probabilities $\pi_{a}$, and the outcomes models $\mu_{a}$ – collectively denoted by $\eta=\{e(\cdot),\pi_{0}(\cdot),\pi_{1}(\cdot),\mu_{0}(\cdot),\mu_{1}(\cdot)\}$). This property occurs because the errors of influence-function-based estimators are products of the errors of nuisance function estimates (call them ${\widehat{\eta}}=\{{\widehat{e}}(\cdot),{\widehat{\pi}}_{0}(\cdot),{\widehat{\pi}}_{1}(\cdot),{\widehat{\mu}}_{0}(\cdot),{\widehat{\mu}}_{1}(\cdot)\}$). Consequently, the overall estimator can achieve a faster rate of convergence even when the individual components of $^{\widehat{\eta}}$ converge at slower rates (see Kennedy (2022) for more details). As a result, we require minimal assumptions on the data distribution to estimate the bounds.
All proposed bounds (and the point identification) can be expressed as linear combinations of elements of the vector $\boldsymbol{\theta}=\mathbb{E}\{\zeta(\mathbf{X})\},$ where $\zeta(\mathbf{x})=\{\mu_{0}(\mathbf{x}),\mu_{1}(\mathbf{x}),\pi_{0}(\mathbf{x}),\pi_{1}(\mathbf{x}),\mu_{0}(\mathbf{x})\pi_{0}(\mathbf{x}),\mu_{1}(\mathbf{x})\pi_{1}(\mathbf{x})\}$. Because the influence function of a linear combination of functionals is a linear combination of the same influence functions Kennedy (2022), it suffices to state the influence function of the components of $\boldsymbol{\theta}$. All bounding quantities can be estimated based on the following influence functions:
@@
The uncentered influence functions for the functionals $\mathbb{E}\{\mu_{a}(\mathbf{X})\}$, $\mathbb{E}\{\pi_{a}(\mathbf{X})\}$, and $\mathbb{E}\{\mu_{a}(\mathbf{X})\pi_{a}(\mathbf{X})\}$ are given by $\varphi_{1,a}({\bf O};\eta),\varphi_{2,a}({\bf O};\eta)$, $\varphi_{3,a}({\bf O};\eta)$, respectively, where
@@
$\varphi_{3,a}({\bf O};\eta)$ $=\varphi_{1,a}({\bf O};\eta)\pi_{a}(\mathbf{X})+\varphi_{2,a}({\bf O};\eta)\mu_{a}(\mathbf{X})-\mu_{a}(\mathbf{X})\pi_{a}(\mathbf{X}).$
@@
Suppose the target of inference is $\theta_{\bf c}={\bf c}^{{\sf\scriptscriptstyle{T}}}\boldsymbol{\theta}$, a linear combination of $\boldsymbol{\theta}$ (and where the elements of $\bf c$} are finite). Collect the influence functions into the vector $\mathbf{\varphi}({\bf O},\eta)=\left\{\varphi_{1,0}({\bf O},\eta),\varphi_{1,1}({\bf O},\eta),\varphi_{2,0}({\bf O},\eta),\varphi_{2,1}({\bf O},\eta),\varphi_{3,0}({\bf O},\eta),\varphi_{3,1}({\bf O},\eta)\right\}.$ All the quantities discussed above, including the bounds, can be written in this form. Corollary 2 in Appendix 9 provides explicit expressions for all previously considered quantities. We now describe the asymptotic behavior of the estimators of these quantities.
@@
${\widehat{\theta}}_{\bf c}-\theta_{\bf c}$ $=\mathbb{P}_{n}[{\bf c}^{{\sf\scriptscriptstyle{T}}}\varphi({\bf O};\eta)-\theta_{\bf c}]+o_{p}(n^{-1/2})+\mathcal{O}_{p}(R_{2}({\widehat{\theta}}_{\bf c})).$
where
$|R_{2}({\widehat{\theta}}_{\bf c})|$ $\lesssim\sum_{a=0,1}\left(\|\mu_{a}(\mathbf{X})-\hat{\mu}_{a}(\mathbf{X})\|\left\{\|\pi_{a}(\mathbf{X})-\hat{\pi}_{a}(\mathbf{X})\|+\|e(\mathbf{X})-\hat{e}(\mathbf{X})\|\right\}\right$.
$\left.+\|e(\mathbf{X})-\hat{e}(\mathbf{X})\|\|\pi_{a}(\mathbf{X})-\hat{\pi}_{a}(\mathbf{X})\|\right).$
Therefore, if $R_{2}({\widehat{\theta}}_{\bf c})=o_{p}(n^{-1/2})$, it follows by the Central Limit Theorem that
$\sqrt{n}\bigl({\widehat{\theta}}_{\bf c}-\theta_{\bf c}\bigr)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{N}\bigl(0,{\bf c}^{{\sf\scriptscriptstyle{T}}}\mathbb{V}\{\varphi({\bf O};\eta)\}{\bf c}\bigr).$
Theorem 1 states that any linear combination of the estimated influence functions of $\boldsymbol{\theta}$ has a limiting normal distribution with asymptotic variance given by the variance of the corresponding linear combination of the true influence functions. This result holds whenever the nuisance estimators converge sufficiently fast so that the second-order remainder term $R_{2}({\widehat{\theta}}_{\bf c})$ is $o_{p}(n^{-1/2})$. A sufficient condition is that the nuisance estimators converge faster than $n^{-1/4}$ in $L_{2}(\mathbb{P})$ norm. Many off-the-shelf machine learning algorithms can achieve these rates under minimal assumptions on the data-generating mechanism Kennedy (2022). An asymptotically valid $(1-\alpha)\times 100\%$ confidence interval may then be constructed as
${\widehat{\theta}}_{\bf c}\pm z_{1-\alpha/2}\sqrt{{\bf c}^{\top}\widehat{\mathbb{V}}\{\varphi({\bf O};\hat{\eta})\}{\bf c}/n},$ (5)
where $z_{q}$ is the $q$-th quantile of the standard normal distribution.
To illustrate, the point identification from Proposition 5 corresponds to choosing ${\bf c}_{1}=(-1,1,0,0,(1-\tau)\delta_{0},(\tau-1)\delta_{1})^{\top}$ so that $\theta_{\bf c}=\Psi_{0}$ under Assumptions 1-5, 6.a, and 7.d. The influence-function based estimator for this quantity is then,
@@
with the corresponding confidence interval given by (5), substituting ${\bf c}_{1}$ for $\bf c$}.
@@
The bound on the second-order remainder term $|R_{2}({\widehat{\theta}}_{\bf c})|$ given in Theorem 1 may involve more error products than are strictly required for specific bounds. The precise second-order remainder terms for individual estimators of each bound are given in Corollary 3 in Appendix 9.
@@
Assuming $^{\widehat{\eta}}$ comes from an independent sample facilitates the statement of Theorem 1 but is not required, provided that cross-fitting is used to separate the estimation of nuisance functions from their evaluation in the influence function (see Chernozhukov et al. (2018)).
Remark 6.
In cases where analysts wish to evaluate the bounds across a grid or range of sensitivity parameters, one may construct uniform confidence bands (that is, confidence intervals with simultaneous coverage guarantees for all parameters in some parameter space) at level $1-\alpha$ using the multiplier bootstrap under minimal additional conditions. We formalize this in Theorem 5 in Appendix 9.3.
Remark 7.
As noted previously, all bounds and identifying expressions may be inverted to obtain tipping points, that is, values of the sensitivity parameters that set the bound or identifying expression equal to zero. In principle, such tipping points may be treated as functionals of the observed data distribution and estimated directly using influence-function based methods. In many cases, these quantities are ratios of linear functionals of $\boldsymbol{\theta}$ so that their efficient influence functions follow from standard results Kennedy (2022). We do not pursue such direct estimators here. Instead, because many analysts will evaluate the bounds over a grid of sensitivity parameters, we recommend inferring tipping points from the sensitivity curve itself, namely as the parameter value at which the lower or upper bound crosses zero. Combined with the uniform confidence bands described in Remark 6, this yields a simultaneous uncertainty quantification for the entire sensitivity curve and therefore an uncertainty region for the tipping point.
@@ 5 Alternative target estimands
In some cases, however, this composite outcome may not be of interest. Building on ideas from the mediation literature Robins et al. (2022), Stensrud et al. (2022) proposed a separable direct effect (SDE) that essentially isolates the effect of the treatment on $Y$. Specifically, the SDE conceives of the treatment indicator $A$ as being comprised of two separate treatments, $A_{y}$ and $A_{d}$, where $A_{y}$ only affects the outcome and $A_{d}$ only affects $U_{I}$. That is, we can conceive of a hypothetical trial where $A_{y}$ and $A_{d}$ are jointly randomized and recover causal contrasts involving the potential outcomes $Y(a_{y},a_{d})$ for different realizations of $a_{y},a_{d}$. Here we are interested in the part of $A$ that affects $Y$ (i.e., $A_{y}$), while holding the part of $A$ that affects $U_{I}$ constant at $A_{d}=0$. We therefore first assume that $Y(a)=Y(a_{y},a_{d})$, and that $U_{I}(a)=U_{I}(a_{y},a_{d})$, and define the SDE as $\Psi_{2}=\mathbb{E}\{Y(1,0)-Y(0,0)\}$. Again, this estimand represents the causal effect via the pathways from $A$ that affect $Y$ alone ($A\boldsymbol{\rightarrow}A_{y}\rightarrow Y$, where the bold arrow indicates a deterministic relationship), leaving the pathways that affect $U_{I}$ unchanged ($A\boldsymbol{\rightarrow}A_{d}\to U_{I}\to Y$). Neither $\Psi_{1}$ or $\Psi_{2}$ are identified under the basic assumptions (Assumptions 1-5). However, these assumptions are sufficient to rewrite $\Psi_{1}$ similarly to how $\Psi_{0}$ was rewritten in Proposition 1.
@@
To bound $\Psi_{2}$, we further require the so-called “dismissible component conditions” Stensrud et al. (2022).
@@
This assumption is required because we do not observe the hypothetical trial that jointly randomizes $(A_{y},A_{d})$. Instead, our observed data contain only the composite treatment indicator $A$, so that $A_{y}=A_{d}$ for all units. This assumption also intuitively formalizes the idea that $A_{d}$ does not affect $Y(a_{y},a_{d})$ except via its effect on $U_{I}$, and similarly, that $A_{y}$ does not affect $U_{I}$. We refer to Stensrud et al. (2022) for more details.
@@
Unlike $\Psi_{0}$, these estimands do not depend on the unobserved quantity $\mu_{a}^{\star}(\mathbf{X})$, and therefore Assumptions 7.b-7.c are not required to bound these quantities. Instead, $\Psi_{1}$ only depends on the unobserved quantities $\pi^{\star}_{a}(\mathbf{X})$ for $a=0,1$, and $\Psi_{2}$ only on $\pi^{\star}_{0}(\mathbf{X})$. In the next section, we provide bounds for these alternative estimands analogous to the bounds in Propositions 4 and 5.
@@ 5.1 Bounds
$\ell_{2}=\tilde{\Psi}-\mathbb{E}\left[\delta_{u0}\pi_{0}(\mathbf{X})(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))\mathbb{I}\{\mu_{1}(\mathbf{X})>\mu_{0}(\mathbf{X})\}\right],$
$u_{2}=\tilde{\Psi}-\mathbb{E}\left[\delta_{u0}\pi_{0}(\mathbf{X})(\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X}))\mathbb{I}\{\mu_{1}(\mathbf{X})\leq\mu_{0}(\mathbf{X})\}\right].$
To illustrate, consider again the running example introduced in Section 3. In this case, $\Psi_{1}\approx 0.17$ and $\Psi_{2}\approx 0.12$. Setting $\delta_{1u}=\delta_{0u}=1$ and $\delta_{1\ell}=\delta_{0\ell}=0$ (which hold by definition), we derive the bounds [0.00,0.33][-0.00,0.33] for $\Psi_{1}$ and $[0.11,0.135]$ for $\Psi_{2}$. These can be further refined using Assumption 6.b.
@@
As noted previously, Assumption 6.a is quite strong, so it is most useful as part of sensitivity analyses, as discussed in Section 3.4. Moreover, for $\Psi_{1}$, once we are simply obtaining point estimates under a range of $\delta_{1},\delta_{0}$, this is essentially equivalent to obtaining bounds $\ell_{1},u_{1}$ under different choices of $\delta_{au},\delta_{a\ell}$.